# Between a Rock and a Hard Place: Exploiting Ethical Reasoning to Jailbreak LLMs

### v2

@@ 5.1 Experimental Setup
Method Turns Llama-3.1-8B Vicuna-13B DeepSeek-V3 GPT-3.5 GPT-4 GPT-4o GLM-4-Plus Claude-3.7-Sonnet
GCG Single 17% 48% 9% 34% 3% 4% 18% 0%
PAP Single 26% 16% 6% 16% 2% 2% 14% 3%
PAIR Single 11% 34% 25% 30% 14% 42% 34% 2%
DRA Single 42% 8% 51% 23% 17% 5% 62% 0%
DeepInception Single 5% 0% 19% 2% 1% 11% 13% 0%
ActorAttack Multi 31% 50% 11% 47% 26% 38% 44% -
Jigsaw Multi 25% 44% 38% 40% 24% 2% 77% 1%
TRIAL (ours) Multi 72% 54% 82% 43% 36% 73% 81% 16%
Table 1: Jailbreak success rates for baseline comparisons and TRIAL on the JBB-Behaviors dataset. The jailbreak success rate was calculated as the total number of successful jailbreaks divided by the total number of prompts, where higher values indicate stronger jailbreak effectiveness. All evaluations were conducted manually in accordance with the 2023 Trojan Detection Challenge (TDC) guidelines to ensure result consistency and accuracy.
Jailbreak Evaluation. While prior jailbreak studies used an LLM-as-a-judge approach [Li et al.(2024a)Li, Dong, Chen, Su, Zhou, Ai, Ye, and Liu], we found it often provides inconsistent results. This gets further complicated by differing LLM judges across benchmarks. To ensure fair evaluation, we adopted a multi-judge strategy: each benchmark’s recommended LLM judge was applied and then supplemented with human manual validation. Specifically, we used Llama-3-70B-Instruct for JBB-Behaviors and AdvBench; the fine-tuned LLaMA-2-13B-cls judge for the HarmBench dataset; and a GPT-4o judge with a scoring-based metric template from the CLAS 2024 Competition for the CLAS dataset. This LLM-based evaluation was further verified through internal manual validation following TDC guidelines [Mazeika et al.(2023)Mazeika, Zou, Mu, Phan, Wang, Yu, Khoja, Jiang, O’Gara, Sakhaee, Xiang, Rajabi, Hendrycks, Poovendran, Li, and Forsyth]. Scoring metrics, classification prompt templates (Tables A5, A6, and A7 in Appendix C), and full manual evaluation details are provided in Appendix C.
Experimental Details. We selected GLM-4-plus as our primary attack model from candidates including Llama-3.1-8B and GPT-4o. GLM-4-plus proved most effective for generating persuasive, adaptive attack queries, whereas more strongly aligned models occasionally refused to generate them. It is also noted that safety alignments of a model can be compromised through methods like fine-tuning with harmful examples [Yi et al.(2024a)Yi, Ye, Chen, Zhu, Chen, Lian, Sun, Xie, and Wu, Qi et al.(2023)Qi, Zeng, Xie, Chen, Jia, Mittal, and Henderson, Yang et al.(2023)Yang, Wang, Zhang, Petzold, Wang, Zhao, and Lin]. However, pursuing this computationally intensive path was beyond the scope of our work. Instead, GLM-4-Plus showcased strong conversational capabilities while being able to generate adversarial queries. This choice yielded high jailbreak success rates. GPT-4o was used for initial clue extraction and scenario generation, with GLM-4-Plus as a backup for this setup phase if GPT-4o refused. Hyperparameter details are in Appendix E.
5.2 Results
Based on Table 1, TRIAL demonstrated strong jailbreak effectiveness by outperforming representative single-turn attacks and showing superior performance against other multi-turn methods across most baseline models. Despite its broad effectiveness, TRIAL encountered greater resistance from Claude-3.7-Sonnet, suggesting its more robust safety alignment. Conversely, TRIAL has high ASR on other newer models generally known for better reasoning capabilities, such as GPT-4o, DeepSeek-V3, and GLM-4-Plus. Our results support the hypothesis that sophisticated reasoning can paradoxically become an exploitable attack vector. To specifically explore the impact of models reputed for deep, human-like reasoning, we conducted a targeted experiment with TRIAL against DeepSeek-R1 [Guo et al.(2025b)Guo, Yang, Zhang, Song, Zhang, Xu, Zhu, Ma, Wang, Bi et al.]. In this evaluation, TRIAL achieved a high ASR of 75%, which reinforced the observation that even advanced or specialized reasoning capabilities do not inherently confer immunity to this style of ethical reasoning exploitation. TRIAL’s strong jailbreaking performance is further showcased across other benchmarks, with detailed results in Appendix D (Tables A8,  A9, and  A10).
6 Evaluating TRIAL against Defense Mechanisms
Defense Attack Llama-3.1-8B DeepSeek-V3 GPT-4o GLM-4-plus
None TRIAL 36% 76% 56% 72%
ActorAttack 36% 52% 40% 44%
LlamaGuard3 TRIAL 12% (–24%) 16% (–60%) 36% (–20%) 20% (–52%)
ActorAttack 28% (–8%) 36% (–16%) 20% (–20%) 24% (–20%)
SmoothLLM TRIAL 36% (0%) 64% (–12%) 20% (–4%) 68% (–4%)
ActorAttack 36% (0%) 44% (–8%) 36% (–4%) 44% (0%)
Circuit Breaker TRIAL 0% (–56%) N/A N/A N/A
ActorAttack 0% (–67%) N/A N/A N/A
Table 2: Success rates of TRIAL and ActorAttack under different defenses. Percentage-point drops (relative to the “None” defense) are highlighted in red. Due to computational constraints, these defense evaluations were performed on 25 evenly sampled instances from the JBB-Behaviors dataset. Results for Circuit Breaker defense were obtained using a Llama3-8B model variant. N/A indicates evaluations were not performed for those models with Circuit Breaker.
Table 2 presents TRIAL’s performance against LlamaGuard3 [Inan et al.(2023)Inan, Upasani, Chi, Rungta, Iyer, Mao, Tontchev, Hu, Fuller, Testuggine et al.], SmoothLLM [Robey et al.(2023)Robey, Wong, Hassani, and Pappas], and Circuit Breaker [Zou et al.(2024)Zou, Phan, Wang, Duenas, Lin, Andriushchenko, Kolter, Fredrikson, and Hendrycks] using ActorAttack as a baseline. LlamaGuard3 significantly curbed TRIAL by detecting harmful intent progressively built via its dynamic, multi-turn ethical framing. This contrasts sharply with ActorAttack’s more predictable, template-driven attacks, which lack the evolving threat signature effectively caught by LlamaGuard3. However, LlamaGuard3 faces challenges, including possible misclassifications and notable computational overhead and latency [Hu et al.(2024)Hu, Piet, Zhao, Jiao, and Wagner]. Conversely, SmoothLLM’s prompt perturbation offered weak resistance with minimal impact on TRIAL. Such defenses are less effective, as TRIAL exploits ethical reasoning via structured dilemmas, rather than specific phrasing vulnerabilities. Although input-output filters can mitigate TRIAL, its dynamic exploitation of ethical reasoning remains a potent challenge for current safeguards. Further, we tested against the Circuit Breaker defense [Zou et al.(2024)Zou, Phan, Wang, Duenas, Lin, Andriushchenko, Kolter, Fredrikson, and Hendrycks], with Llama3-8B. Both baselines attack success rate dropped to 0%. Notably, this defended Llama3-8B resisted TRIAL’s initial ethical dilemma and pull back query entirely, suggesting such defenses can preempt TRIAL’s manipulative framing at an early stage.
6.1 Ablation Study: Longer Attack Sequences on Resistant Prompts
To assess the impact of prolonged engagement, we extended TRIAL interactions on JBB-Behaviors prompts initially resisted by GPT-4o. Extending interaction length ($K$) increased overall jailbreak success for these challenging prompts (Figure 3a), though conversions primarily occurred in earlier extended rounds with diminishing returns. When prompts remained uncompromised, GPT-4o deployed diverse refusal strategies (Figure 3b,  3c). While extending $K$ can overcome some initial refusals, its effectiveness ultimately plateaus. Two primary factors limit continued success: (1), prolonged discussions risk derailment if LLMs misinterpret increasingly nuanced attacker queries; (2), strongly aligned models inherently resist sustained engagement on the most challenging prompts. In these later stages, such models are adept at issuing hard rejects to even subtly or indirectly phrased harmful requests, and attacker queries that become overly implicit or unclear can also independently lead to refusals.
(a) Cumulative conversion over extended rounds.
(b) Final outcomes after 8 extended rounds.
(c) Evolution of resistance strategies.
Figure 3: Extending TRIAL interactions (up to K=10 rounds) for 30 JBB-Behaviors prompts initially failing against GPT-4o (victim; GLM-4-plus attack model). The first two extended rounds were not evaluated for jailbreak success, affecting cumulative data in (a). Left: (a) Cumulative jailbreak success for these prompts on GPT-4o over extended rounds. (b) Final outcomes for the 30 prompts after extension, human-labeled: ’Successfully Jailbroken’ (harmful content generated); ’Hard Reject’ (explicit refusal); ’Soft Reject’ (indirect refusal/disengagement); ’Compliant Discussion’ (non-harmful engagement); ’Evasion/Off-Topic’ (avoidance/irrelevant response). Right (c): Evolution of GPT-4o’s resistance strategies during extended interaction, using categories from (b).
6.2 Qualitative Evaluation
We observed that the JailbreakBench judge occasionally misclassified jailbreak scores. In many cases, although victim responses provided vague and basic instructions that contained harmful elements relevant to the original prompt, they failed to fully execute the requested task, such as generating code or composing articles/emails. This issue was particularly evident in smaller models, such as Vicuna and Llama, which struggled to produce complete outputs despite demonstrating partial compliance. Furthermore, the jailbreak judge’s classification appears to exhibit a false-negative bias when evaluating responses, particularly responses that include caveats or warnings.
To illustrate the effectiveness of TRIAL, we provide examples across various attack categories, as shown in Appendix F. Additionally, some chat histories have been truncated to balance toxicity while preserving the integrity of the qualitative analysis.
6.3 Empirical Analysis
LLMs exhibit exploitable ethical inconsistencies in forced dilemmas. Our empirical analysis, supported by Figure A2 (Appendix E, showed that while models frequently selected the harmful Option A, they sometimes chose Option B even if it appeared to entail greater negative consequences (e.g., sacrificing more people). This suggests models can exhibit inconsistent or incomplete ethical reasoning when weighing presented dilemmas. Crucially, an initial selection of Option B did not prevent jailbreak; TRIAL’s adaptive querying and ’pull back query’ designed to re-engage the victim model, often successfully nudged it to reconsider and select Option A, especially when the dilemma framed options with closely balanced perceived negative utility. However, the effectiveness of this simple re-prompting was limited and consistently struggled when Option A embodied universally condemned harms, if the scenario failed to convincingly frame Option A as the ’lesser evil’. These observations clearly show the pull-back query’s limits. While it can reverse weaker initial refusals, it consistently fails when models enforce robust ethical principles or safety protocols. Therefore, future work must develop more adaptive re-engagement strategies for highly resistant scenarios.
7 Conclusion and Discussion
We introduced TRIAL, a novel multi-turn jailbreak framework that successfully exploits LLM ethical reasoning using trolley problem-inspired dilemmas to bypass safeguards on state-of-the-art models. Our work reveals that LLMs’ dynamic reasoning processes are vulnerable. Iterative, context-aware prompts can steer models toward compliance even after initial refusals. This underscores the perceived relativity of harm as a significant attack vector, though TRIAL’s effectiveness may be limited in scenarios involving morally absolute issues where no “lesser evil" justification is plausible (e.g. sexual exploitation of minors).
@@ 7 Conclusion and Discussion
While TRIAL successfully navigates the ethical reasoning of many models, our evaluation against the Circuit Breaker defense on Llama3-8B revealed its profound impact. We also observed that this heavily defended model occasionally issued hard rejects even at the seemingly benign stage of choosing an option within the ethical dilemma. This suggests that some highly robust defenses might also exhibit over-sensitivity, potentially penalizing or refusing engagement with multi-step interactions even if the initial steps are not inherently harmful. Such behavior underscores the ongoing challenge in developing defenses that are both resilient to sophisticated attacks like TRIAL and permit benign interactions. A critical consideration for future AI safety research is striking an effective balance between ensuring model harmlessness and avoiding over-refusal tendencies that can hinder legitimate uses and the model’s overall utility.
Our empirical analysis further underscored that LLMs do not consistently minimize harm in ethical dilemmas and can be persuaded by TRIAL’s "lesser evil" framing to reverse safer initial stances. This reveals that LLM security is highly context-dependent, with models susceptible to justificatory framing that erodes safeguards over successive interactions, especially when faced with complex ethical trade-offs. This dependency extends not only to the victim model’s characteristics but also critically to the attack model itself. Although our experiments used a weakly aligned model (GLM-4-Plus, as detailed under Section 5.1) for attack generation, the effectiveness of TRIAL could be amplified if the attack model were specifically fine-tuned for adversarial purposes [Yi et al.(2024a)Yi, Ye, Chen, Zhu, Chen, Lian, Sun, Xie, and Wu, Qi et al.(2023)Qi, Zeng, Xie, Chen, Jia, Mittal, and Henderson, Yang et al.(2023)Yang, Wang, Zhang, Petzold, Wang, Zhao, and Lin]. Such a fine-tuned attacker might exhibit fewer inhibitions in generating harmful queries and could craft even more persuasive attack prompts and potentially overcoming more robust defenses. This potential for escalating attacker sophistication underscores an additional dimension of vulnerability. Consequently, these findings urgently call for the development of dynamic, context-aware defense mechanisms capable of recognizing and interrupting such multi-turn adversarial reasoning.
Limitations
TRIAL’s primary methodological limitation is its dependence on framing requests as trolley problem-like dilemmas, requiring the harmful action (Option A) to be a ’lesser evil’ than a catastrophic alternative (Option B). This restricts its adaptability for prompts not fitting this utilitarian trade-off, especially ’morally absolute’ issues (e.g., generating content on child exploitation). For such issues, the profound intrinsic harm thwarts any ’greater good’ justification an LLM might accept, causing the ’lesser evil’ framing to fail as ethical weight overrides utilitarian arguments. Thus, while potent where harms can be ethically weighed, TRIAL is not universally applicable and struggles with actions deemed unjustifiable regardless of context
Furthermore, an operational constraint can limit our approach when the chosen attack or helper models are themselves too stringently safety-aligned. These models may refuse to generate the initial trolley problem scenario or the subsequent persuasive attack queries if they detect the underlying harmful intent. Such refusals can prevent the TRIAL process from commencing effectively, thereby limiting its practical application depending on the models employed.
Ethics Statement
This paper presents a systematic jailbreaking attack that enables users to generate harmful content from publicly released LLMs. A study has mentioned that LLMs, despite their impressive performance on ethical reasoning benchmarks, do not truly understand ethical principles but instead rely on statistical correlations and surface-level lexical features [Albrecht et al.(2022)Albrecht, Kitanidis, and Fetterman]. Hence, it is compulsory to develop a deeper understanding of how models reason through various dilemmas involving ethical decision-making to mitigate this issue. We strictly adhered to ethical guidelines throughout our research.
Through a comparison of various LLMs, we found that TRIAL performs exceptionally well on the latest models compared to older ones. This observation shows one way to mitigate our attack. However, as LLMs continue to improve in reasoning, our research aims to highlight the paradox between enhanced reasoning and emerging safety vulnerabilities. We hope that our findings will contribute to a better understanding of this trade-off and inform the development of more robust and secure AI systems.
In summary, this research is conducted with the aim to improve LLM safety, not for any malicious usage. We also hope that our work will provide insight towards future research towards safety alignment, particularly in the direction of ethical decision making. We commit to ongoing monitoring and updating our research in line with technological advancements.
Appendix A Detailed TRIAL Algorithm
Input: Harmful query $\zeta$, attack model AA, victim model $V$, judge model $J$, helper model $H$, iterations $k$,
Output: Victim responses, $r_{i}$
Initialize: Conversation history $H_{C}\leftarrow[\,]$
$C_{\zeta}\leftarrow\text{generate\_clue}(\zeta,H)$
$S_{\zeta}\leftarrow\text{generate\_scenario}(C_{\zeta},H)$
for $i=1$ to $K$ do
  if $i==0$ then
     $q_{i}\leftarrow\text{init\_attack}(S)$ ;
     $r_{i}\leftarrow\text{get\_victim\_response}(H_{c},V)$ ;
     if $r_{i}$ contains Option B then
        $q_{i}\leftarrow\text{pull\_back}()$; # Change to Option A
        $H_{c}\leftarrow H_{C}+{\{q_{i},r_{i}\}}$
        continue
 $q_{i}\leftarrow\text{attack}(H_{C},C_{\zeta},A)$
  $r_{i}\leftarrow\text{get\_victim\_response}(H_{C},V)$ ;
  $s\leftarrow\text{judge}(x,r_{i},J)$
  $H_{C}\leftarrow H_{C}+{\{q_{i},r_{i},s\}}$
  if $s==\text{True}$ then
     return $H_{c}$
return $H_{c}$
Algorithm 1 TRIAL: Multi-turn Jailbreak
Appendix B Difference Between Single and Multi-turn Jailbreak Attacks
Figure A1 demonstrates a main difference between single and multi-turn jailbreak techniques. In a single-turn attack, the original harmful prompt is directly transformed into an adversarial prompt using token-based or prompt-based methods. In contrast, a multi-turn jailbreak attack decomposes the original harmful prompt into smaller subprompts. Multi-turn jailbreaking process involves iterative refinement of the adversarial query over multiple dialogue turns.
Figure A1: An example of single-turn and multi-turn jailbreak attack
Appendix C TRIAL Chat Templates

### v3

@@ D.1 Additional Evaluation on TRIAL
Method Turns Llama-3.1-8B DeepSeek-V3 GPT-4o GLM-4-plus
GCG Single 34% 14%
PAP Single 30% 11% 23% 22%
ActorAttack Multi 70% 60% 81% 73%
Jigsaw Multi 54% 68% 6% 64%
TRIAL (ours) Multi 74% 86% 86 % 84%
Open-Source Closed-Source
Method Turns Llama-3.1-8B DeepSeek-V3 GPT-4o GLM-4-plus
GCG Single 14% 3%
PAP Single 55% 39% 43% 41%
ActorAttack Multi 44% 58% 42% 62%
Jigsaw Multi 64% 53% 12% 42%
TRIAL (ours) Multi 74% 86% 86 % 84%
Open-Source Closed-Source
Method Turns Llama-3.1-8B Qwen-2.5-7B GPT-4o GLM-4-plus
GCG Single 27.00 33.00 18.75 6.25
PAP Single 17.00 57.25 66.25 56.25
ActorAttack Multi 30.00 71.25 37.50 35.00
Jigsaw Multi 66.25 42.50 56.00 6.25
TRIAL (ours) Multi 56.25 91.25 95.00 87.50
ASR ($\downarrow$)
Defense Attack Llama DS-V3 GPT-4o GLM-4
None TRIAL 36 76 56 72
ActorAttack 36 52 40 44
Self Reminder TRIAL 32-4 72-4 24-32 56-16
ActorAttack 33-3 46-6 16-24 41-3
LlamaGuard3 TRIAL 12-24 16-60 36-20 20-52
ActorAttack 28-8 36-16 20-20 24-20
SmoothLLM TRIAL 360 64-12 52-4 68-4
ActorAttack 360 44-8 36-4 440
Table A4: Defense evaluation on TRIAL and ActorAttack. Subscripts denote absolute percentage-point change relative to the None defense for the same attack and model. This experiment is evaluated on sampled 50 JBB-Behaviors instances.
ASR ($\downarrow$) FPR ($\downarrow$) ACC ($\uparrow$)
Models None PAP TRIAL XsTest FalseReject MMLU
Kimi-K2-Thinking 0% 1% 6% 0% 0% 76%
DeepSeek-R1-0528 6% 6% 32% 5% 8% 84%
Qwen3-32B 3% 23% 46% 2% 5% 82%
ERR-LlaMA 0% 0% 0% 14% 60% 56%
ERR-Mistral 0% 0% 3% 6% 14% 56%
Table A5: Comparison with open-source large reasoning models. The same StrongReject dataset is used for evaluating ASR. Bold represents the best results.
(a) GCG
(b) PAP
(c) DRA
(d) PAIR
@@
Appendix E Experimental Setup Details
E.1 Jailbreak Attack Baselines
Greedy Coordinate Gradient (GCG)
We followed the original setting of GCG Zou et al. (2023b), replacing the loss function with the mellowmax loss Asadi and Littman (2017) to improve performance. This adjustment maintained efficiency while leveraging the benefits of the mellowmax loss function. Additionally, we did not enforce early stopping, ensuring that the optimization process ran for the full number of steps. Due to resource constraints, we employ transfer attacks on black-box models and DeepSeek-V3.
Persuasive Adversarial Prompt (PAP)
Prompt Automatic Iterative Refinement (PAIR)
ActorAttack
Jigsaw Puzzles
DRA
The Disguise and Reconstruction Attack (DRA) is a black-box jailbreak method for large language models (LLMs) that exploits biases in safety fine-tuning by disguising harmful instructions within queries to evade direct rejection, then prompting the model to reconstruct the original harmful instruction in its completion through payload reconstruction and context manipulation, effectively bypassing internal safeguards and inducing unethical responses. For replication, we use the default settings mentioned in the paper, including the use of base64 encoding for disguise, no reliance on additional LLMs for prompt optimization, and evaluation on models like GPT-4 with minimal query trials.
DeepInception
DeepInception is a lightweight, training-free jailbreak approach inspired by the Milgram shock experiment, which hypnotizes LLMs by injecting nested imaginary scenes that leverage the model’s personification abilities to create a self-losing state under authority, allowing adaptive overriding of safety guardrails through jointly and continually inducing harmful content in a virtual multi-layer fiction. For replication, we use the default settings mentioned in the paper, such as the universal prompt template with recursive nested instructions. We optimized the prompts using Llama3.1-8B-Instruct.
E.2 Safety Alignment Baseline
To strictly evaluate the robustness of our defense, we compare our method against three state-of-the-art safety alignment baselines. We utilize the open-sourced checkpoints available on Hugging Face, specifically focusing on variants based on Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.2.
Circuit Breakers
Zou et al. (2024) Proposed as a mechanism to mitigate adversarial attacks, Circuit Breakers employ a representation engineering approach. Instead of relying on standard refusal training, this method identifies harmful directions within the model’s internal representations and trains a “circuit breaker” to actively reroute these activations into incoherent states, thereby short-circuiting the generation of harmful content.
RepBend
Yousefpour et al. (2025) Representation Bending (REPBEND) introduces a fine-tuning framework designed to geometrically separate the latent representations of harmful and benign queries. By optimizing a loss function that enforces orthogonality between safe and unsafe activation trajectories, REPBEND aims to make the model’s refusal behavior robust against jailbreak attempts while preserving general utility.
RATIONAL
Zhang et al. (2025a) Moving beyond binary refusal, RATIONAL (Reasoning-Enhanced Fine-Tuning) trains models to generate an explicit safety rationale before producing a final response. This approach leverages the model’s reasoning capabilities to analyze the intent and ethical implications of a query, aiming to provide context-aware safety judgments rather than heuristic-based rejections.
E.3 Jailbreak Benchmarks
This section provides an overview of the datasets we used from our jailbreak experiments.
JailbreakBench: JBB-Behaviors Dataset
CLAS 2024: Jailbreaking Attack Track (JAT) Datasest
HarmBench Behavior Dataset
HarmBench is another standardized evaluation framework for automated red teaming, where they provide both a dataset of harmful prompts and a corresponding text classifier. We selected GLM-4-plus as the attack model, while Llama-2-13b-cls, a fine-tuned version of Llama 2 by Harmbench, was chosen as the judge model. Due to the large size of this dataset, we sampled it to 100 prompts which has even distribution across these categories.
AdvBench Dataset
AdvBench comprises 500 instruction‐style prompts representing a broad spectrum of malicious or disallowed behaviors. Unlike the harmful string setting, where each input is evaluated independently, the attacker’s objective is to discover a single adversarial string that will induce the model to produce a compliant (and therefore unsafe) response across as many of these behaviors as possible. For our experiments, we used the AdvBench subset sampled by Chao et al. (2023).
E.4 Data Curation for Alignment Tuning
We curated balanced harmful and benign scenarios to train our ERR defense. For the Explain Mode, we generated harmful scenarios by drawing harmful prompts from established jailbreak and red-teaming datasets, including JBB-Behaviors, JAT, HarmBench, AdvBench, and LAT. For the Engage Mode, designed to promote helpful responses on safe queries, we curated benign scenarios by sampling 1,500 instructions from the Alpaca-cleaned dataset. This dual-mode curation ensures that the model learns to distinguish harmful from benign contexts while maintaining high helpfulness and reducing over-refusals.
LAT Dataset
The LLM‑LAT (Sheshadri et al., 2024) dataset provide a structured collection for evaluating and improving LLM safety via latent adversarial training. It consists harmful dataset with roughly 5k examples containing malicious or unsafe prompts. These datasets are designed to support both training and evaluation, enabling models to learn to distinguish and appropriately respond to harmful content while preserving general utility. We sample this dataset to 1k examples for our Explain/Engage alignment data.
StrongREJECT Dataset
StrongREJECT (Souly et al., 2024) is a small benchmark of malicious prompts designed for evaluating large language models’ susceptibility to jailbreak attacks. It comprises roughly 313 “forbidden” or harmful prompts that a safely aligned model should refuse or handle appropriately, and is used to assess whether a jailbreak method enables a model to produce harmful outputs rather than refuse them. We sample this dataset to 100 even examples, as it serves as our test set for ERR baseline in Tables 2 and A5.
Alpaca
Alpaca dataset (Taori et al., 2023) contains 51,760 English-language instruction-following examples derived from Stanford’s Alpaca (generated via OpenAI’s text-davinci-003 using Self-Instruct), with cleaned data to fix issues for better quality; it includes fields like instruction, output, and text, covering tasks such as classification, summarization, code generation, and factual queries. This dataset is primarily used for instruction-tuning pretrained LLMs to improve adherence to user instructions in controlled studies. For replication, we sample to 1500 prompts from the dataset, which acts as our benign training data.
E.5 Overrefusal Benchmarks
This section provides an overivew of the overrefusal evaluation benchmarks for ERR baseline. We sample evenly to 100 prompts from the dataset.
XsTest
XsTest (Röttger et al., 2024) is a test suite designed to identify exaggerated safety behaviors in large language models (LLMs), focusing on overrefusal where models excessively reject benign prompts that are not clearly unsafe, while ensuring appropriate compliance with safe prompts and refusal of unsafe ones.
PHTest: Automatic Pseudo-Harmful Prompt Generation for Evaluating False Refusals in Large Language Models
PHTest (An et al., 2025) is an evaluation dataset introduced in the paper which automatically generates diverse, content-controlled, and model-dependent pseudo-harmful prompts that are actually harmless (e.g., "how to kill a mosquito") to assess false refusals in LLMs. This benchmark revealed that many defenses increase false refusal rates and undermine usability.
FalseReject: A Resource for Improving Contextual Safety and Mitigating Over-Refusals in LLMs via Structured Reasoning
FalseReject (Zhang et al., 2025b) is curated by Amazon, comprising 16k seemingly toxic queries with structured responses across 44 safety categories, generated using a graph-informed adversarial multi-agent framework to create diverse prompts and explicit reasoning for distinguishing safe from unsafe contexts. It includes datasets for instruction-tuned and reasoning-oriented models, a human-annotated benchmark, and demonstrates that finetuning with it reduces over-refusals in 29 benchmarked LLMs without compromising safety or capabilities.
E.6 General Benchmarks
This section provides an overview of the helpfulness evaluations benchmarks for ERR baseline. We also sampled all benchmarks to 100 prompts evenly.
MMLU: Measuring Massive Multitask Language Understanding
MMLU (Hendrycks et al., 2021) is a benchmark that evaluates a text model’s multitask accuracy across 57 tasks spanning subjects like elementary mathematics, US history, computer science, and law, aiming to measure extensive world knowledge and problem-solving abilities to assess the breadth and depth of an LLM’s academic and professional understanding.
GSM8K: Grade School Math 8K
GSM8K (Cobbe et al., 2021) is a dataset introduced in the paper "Training Verifiers to Solve Math Word Problems," consisting of 8.5K high-quality, linguistically diverse grade school math word problems to diagnose failures in multi-step mathematical reasoning for LLMs. It proposes training verifiers to rank candidate solutions for improved performance, providing a benchmark to evaluate and enhance LLMs’ reasoning capabilities, thereby contributing to their helpfulness in solving math tasks.
HumanEval
HumanEval (Chen et al., 2021) is a dataset of 164 programming problems in Python, each with a function signature, docstring, body, and unit tests, designed to evaluate LLMs’ code generation abilities by prompting models to produce functional code that passes the tests.
E.7 Evaluation Metrics
We evaluate TRIAL on attack success rate, ERR across harmfulness (drop in ASR), over-refusal, and general helpfulness.
Attack Success Rate (ASR).
[…]
where $y_{i}$ is the model’s response (after ERR intervention), and the indicator function evaluates whether the response is harmful according to our LLM-as-a-Judge classification.
Over-Refusal Rate (ORR).
[…]
where refusal is defined as a response that avoids engaging the benign task (i.e., outputs a refusal token sequence).
General Utility / Task Accuracy.
Assesses performance on standard reasoning or knowledge benchmarks. For instance, on MMLU:
[…]
where $y_{k}^{\star}$ is the ground-truth answer, and $y_{k}$ is the model prediction after ERR intervention.
GPT-4o Judge
Since this judge uses a scoring based evaluation metric, the harmfulness score for each victim response is evaluated using the HEX-Phi scoring system Qi et al. (2024), which assigns scores ranging from 1 to 5 and is mapped to a range of 0 to 100 for Attack Success Rate (ASR) calculation. If no jailbreak response is detected (score = 5), the ASR calculation will select the victim response with the highest rating. The ASR is computed by summing the mapped scores for all victim responses, dividing by the maximum possible score (i.e., 100N100*N, where $N$ is the number of responses), and expressing the result as a percentage.
Manual Evaluation
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
  6. 6.
E.8 Hyperparameters and Training Details
TRIAL Hyperparameters
ERR Training Hyperparameters
We report the recommended hyperparameters based on our experiments, applied consistently across both Llama-3.1-8B-Instruct and Mistral-7B-Instruct-v0.2.
Stage 1: Gate Training.
  • Gate layer: $l_{d}=12$
  • Batch size: $128$
[…]
  • Optimizer: AdamW (weight decay =0.01=0.01)
  • Epochs: 33
  • Sparsity coefficient: $\lambda=0.1$
Stage 2: Adapter Training.
  • Intervention layers: $l_{s}=16$ to $l_{e}=31$
  • LoRA rank: $r=16$, scaling factor: $\alpha=0.5$
  • Safety floor: $\epsilon=0.01$
  • Batch size: $64$
[…]
  • Optimizer: AdamW (weight decay =0.01=0.01)
  • Epochs: 55
  • Gradient clipping: $1.0$
  • Mixed precision: FP16
Hardware.
[…]
Appendix F TRIAL Attack Examples and Their Corresponding Scenarios
We provided four examples of TRIAL attacks using different harmful prompts targeting GPT-4o and Llama-3.1-8B-Instruct in Figure A5 and  A6. The red chat represents the attacker’s prompt, while the blue chat represents the victim’s response. The red dotted line shows the exact harmful information that corresponds to the harmful prompt. We have truncated the chat history to limit page length as well as exclude explicit harmful content and ensure safety, while still preserving the context necessary for quantitative analysis.
Appendix G TRIAL Chat Templates